OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 4x2 + 24x + 32

OpenStudy (firekat97):

the vertex lines on the axes of symmetry for a parabola, so first you want to find that and you can do so by using axes of symmetry = -b/2a where y = ax^2 + bx + c

OpenStudy (firekat97):

lies*

OpenStudy (anonymous):

f(x) = 4x2 + 24x + 32 b=24 a=4 -b/2a -24/8

OpenStudy (firekat97):

yup thats correct! so now to find the y coordinate that goes with this, we simply sub -24/8 = -3, into the equation, to solve for y

OpenStudy (anonymous):

Hmm how did you get the -3 ? Sorry I'm really bad at math !!

OpenStudy (firekat97):

Oh thats okay! I just simplified -24/3 because 3 goes into 24, 8 times and we have a minus out the front as well... @ashhhrodriguez

OpenStudy (anonymous):

Oh WOW thank you! What exactly do I sub 3 for all the x values ?

OpenStudy (anonymous):

I mean -3!

OpenStudy (firekat97):

okay so we know that at our axes of symmetry, x = -3, right?

OpenStudy (firekat97):

Now we want to sub this back into the equation, to find out what y is equal to as x = -3... agreed?

OpenStudy (anonymous):

Ohh so we already have our x value to our vertex -3

OpenStudy (firekat97):

yuppp thats what the equation -b/2a gave us

OpenStudy (firekat97):

so now we have to find the y value which goes with it

OpenStudy (firekat97):

and we can do that by subbing x = -3 back into the original equation

OpenStudy (anonymous):

(-3,-4) Thank you so so much !

OpenStudy (firekat97):

Yup thats correct! And no problem! :)